OpenStudy (anonymous):

Write an equation for the following condition: Perpendicular to y= -3x+7, through (3,5)

OpenStudy (anonymous):

So the equation: y= -3x+7 has a slope of -3. The question is asking for a line that is perpendicular to that line, which means the new slope must be inverse negative So, new slope: 1/3 Use the point (3,5) Now use the point slope formula y-y1=m(x-x1) y-5=1/3(x-3) Now solve for y in the following equation: y-5=1/3(x-3) y=1/3x+4. Therefore the equation for a perpendicular line to y= -3x+7, and that goes through the point (3,5), is : y=1/3x+4.